The sum of the first two digits of a three digit number is 4. The sum of the numbers
that is formed by using each digit of the number once is 1998. If the number has 8
factors, then what is the number?
1
Expert's answer
2020-10-25T18:34:02-0400

Let the number is "\overline{abc}." Then

"a+b=4, a\not=0"


"100a+10b+c+100a+b+10c+"

"+100b+10a+c+100b+a+10c+"

"+100c+10a+b+100c+a+10b=1998"

"222(a+b+c)=1998"

"a+b+c=9"

"c=9-4=5"

The possible numbers are


"135, 225, 315, 405"

The 8 factors of 135 are: "1,3,5,9,15,27,45,135"

The 9 factors of 225 are: "1,3,5,9,15,25,45, 75,225"

The 12 factors of 315 are: "1,3,5, 7,9,15,21, 35, 45, 63,105, 315"

The 10 factors of 405 are: "1,3,5, 9,15,27, 45, 81,1305, 405"


Therefore the number is "135."